# Origin of the Very High Energy -rays in the Low-luminosity Active Galactic Nucleus NGC 4278

### v1

@@ 3 X-ray Emission of the nucleus in NGC 4278
The X-ray light curve of NGC 4278 is compiled in Figure 1, incorporating the archived X-ray data from Younes et al. (2010) and the analysis results from the Chandra and Swift-XRT observations conducted in this study. Notably, NGC 4278 exhibits significant flux variations in the X-ray band. During the observation on MJD 59546, Swift-XRT detected a remarkably high-flux state for NGC 4278 in the X-ray band, surpassing even the historically highest flux obtained from XMM-Newton observation on 2004 May 23. The ratio between the highest flux observed by Swift-XRT in 2021 and the lowest flux observed by Chandra in 2010 exceeds one order of magnitude. Unfortunately, no additional data between 2010 to 2021 could be acquired to estimate variability timescales accurately. While previous studies have reported a 10% increase in flux over a period of 1 hr during XMM-Newton observations (Younes et al. 2010), due to limited exposure time during the Swift-XRT observations, further investigation into short timescale variability is not feasible in this study.
The compact and symmetric two-sided radio morphology of NGC 4278 fulfills the criteria for classification as a CSO, and it is even included in a ‘bona fide’ CSO sample (Kiehlmann et al., 2024). CSOs may be transient or episodic sources (O’Dea & Saikia 2021 for a review), with some -ray emitting CSOs reported to display episodic nuclear activity (Lister et al. 2020; Gan et al. 2021, 2022). It should be noted that radio observations at 6 cm from 1972 to 2003 indicate variability in the flux density of NGC 4278, including an observed outburst around 1985, which may be associated with the presence of an active nucleus (Giroletti et al. 2005). Therefore, the high-flux state observed in X-rays for NGC 4278 could potentially be attributed to the reactivation of its nucleus once again around 2021. This may be the reason that NGC 4278 has recently been detected at the VHE -ray band by LHAASO.
The photon spectral index () against the X-ray flux is depicted in Figure 2. Despite the presence of large errors in the data points, significant spectral variations can still be observed. NGC 4278 shows the harder when brighter behavior, which is commonly observed among BL Lacs in the X-ray band (e.g., Tramacere et al. 2009; Tramacere et al. 2011; Zhang et al. 2013; Di Gesu et al. 2022; Hu et al. 2024).
@@ 4 Constructing and Modeling the SED of Nucleus
The reactivation of the nucleus, as mentioned above, likely result in the detection of VHE -rays and a high-flux state in X-rays from NGC 4278. Consequently, it is plausible that both the emission in X-rays and -rays originate from the same region. In order to further investigate the -ray emission properties of NGC 4278, we collect its multiwavelength observation data from literature and construct a broadband SED. However, only a few non-simultaneous observations from Younes et al. (2010) are available in the radio-optical-UV bands, as shown in Figure 3. The sensitivity curves of LHAASO (Wang et al. 2022), VERITAS (Bigongiari & CTA Consortium 2016), and Fermi-LAT (Funk et al. 2013) are also presented in Figure 3. The hard X-ray spectrum and soft TeV -ray spectrum should correspond to two distinct radiation components, resembling those typical TeV-emitting BL Lacs that have a high synchrotron radiation peak (e.g., Zhang et al. 2012). We employ a simple one-zone leptonic model, widely used in modeling the broadband SEDs of TeV BL Lacs (e.g., Zhang et al. 2012; Yan et al. 2014), to reproduce the SED of NGC 4278, only considering synchrotron and synchrotron-self-Compton (SSC) processes of the relativistic electrons.
@@
The emission region is assumed as a sphere with radius , magnetic field , and a Doppler boosting factor , where is taken as . BL Lacs generally show variability on daily or hour timescales (e.g., Xie et al. 2001; Ravasio et al. 2002; Błażejowski et al. 2005; Fossati et al. 2008; Acciari et al. 2011; Nalewajko 2013), we thus take day for NGC 4278. The VLBA observations indicate that NGC 4278 has mildly relativistic jets () with a viewing angle of , which result in a Doppler booting factor of 2.7 (Giroletti et al. 2005). We thus fix to be 2.7. The values of and are constrained by the spectral indices at X-ray and TeV bands. Since cannot be constrained and is fixed as , while is poorly constrained by the last observation point at TeV band. We adjust the parameters , , and to represent the observed SED of NGC 4278. No detection of VHE -rays from NGC 4278 has been reported by other TeV detectors; therefore, we consider the sensitivity curve of VERITAS as an upper limit during SED modeling. The Klein–Nishina effect (Ghisellini et al. 1998) and the absorption of extragalactic background light (EBL; Finke et al. 2022) are also taken into account in the SED modeling. The SED fitting parameters are given in table 2.
Reproducing the SED of NGC 4278, we obtain G, , and cm-3. It should be noted that the fitting parameters and are degenerate (Zhang et al. 2012). If a high magnetic field strength is required to represent the broadband SED of NGC 4278, then a smaller . For example, would be necessary with G. Based on these SED fitting parameters, we also calculate the jet power () in case of the e± pair jet for NGC 4278 by assuming that the jet consist of electrons (), magnetic fields (), and radiation (). The corresponding values are erg s-1, erg s-1, and erg s-1, respectively (see Table 2).
@@ 5 Discussion
The angular resolution of LHAASO can achieve around 10 TeV (Ma et al. 2022). As stated in Section 2.1, there are two other GeV-TeV BL Lacs located within a distance of less than from NGC 4278. Moreover, NGC 4278 harbors a substantial number of LMXBs (Brassington et al. 2009; D’Abrusco et al. 2014), and -rays have been observed from several LMXBs in the Galaxy. We are wondering whether the VHE -rays detected by LHAASO from 1LHAASO J1219+2915 could potentially originate from either the two -ray emitting BL Lacs or these LMXBs in NGC 4278.
@@ 5.1 Two TeV BL Lacs Located near NGC 4278
The GeV-TeV emitting BL Lacs 1ES 1218+304 and B2 1215+30 are situated at a distance of 0.929 and 0.964, respectively, from the radio position of NGC 4278. Firstly, we conduct a literature search to identify the highest observed flux at the TeV band for the two BL Lacs. We then extrapolate their intrinsic -ray spectrum up to 100 TeV assuming a power-law spectral form. Subsequently, we calculate the observable spectral shape by taking into account the absorption of EBL and compare it with that of 1LHAASO J1219+2915. Two EBL models are considered in this analysis, corresponding to the strongest absorption (Franceschini & Rodighiero 2017) and weakest absorption (Kneiske & Raue 2010), respectively.
@@
The LHAASO is capable of detecting both 1ES 1218+304 and B2 1215+30 when they are in a high-flux state, as depicted in Figures 4. However, the estimated observable spectra of the two BL Lacs fail to represent the observed highest energy photon of 1LHAASO J1219+2915 when considering the strongest EBL absorption model. In other words, due to their high redshift, it appears impossible for the LHAASO to detect -ray photons with energies exceeding 10 TeV from the two BL Lacs, even when considering the weakest EBL absorption model. Therefore, the VHE -ray emission from 1LHAASO J1219+2915 should not be contributed by the two BL Lacs.
@@ 5.2 LMXBs in NGC 4278
Several Galactic LMXBs have been detected at the -ray band (e.g., Xing & Wang 2015; Harvey et al. 2022a; Harvey et al. 2022b; Kantzas et al. 2022). In NGC 4278, there exist multiple LMXBs with an X-ray luminosity exceeding erg s-1 (Brassington et al. 2009; Fabbiano et al. 2010), which are likely to be sources of -ray emission as well. Therefore, we estimate the potential -ray flux from these LMXBs in NGC 4278 by assuming that they exhibit a similar X-ray to -ray flux ratio as the Galactic -ray emitting LMXBs, i.e.,
@@
We extrapolate the GeV spectrum of PSR J1023+0038 up to the TeV band by assuming they follow the same power-law distribution (the black solid line in Figure 5). Multiplying the black solid line by 0.19, we obtain the possible -ray flux (the blue solid line) of the brightest LMXB in NGC 4278. The blue dashed line in Figure 5 presents the result after considering EBL absorption. It is evident that if there is only one -ray emitting LMXB in NGC 4278, its emission would not be detectable by Fermi-LAT and VERITAS but could potentially be detected by LHAASO. NGC 4278 hosts several LMXBs with X-ray luminosities exceeding erg s-1, some of which exhibit significant variability (Fabbiano et al. 2010). Therefore, it is plausible that the observed -rays from 1LHAASO J1219+2915 can be attributed to some LMXBs in NGC 4278. However, no detection has been made for any LMXB at TeV energy band.
In addition, the spectrum at the TeV band of LMXB may not follow the same power-law spectral shape as that at the GeV band; a flatter spectrum is observed in high-mass X-ray binaries (HMXBs) at the TeV band, for example, HESS J1018–589 A/1FGL J1018.6–5856 (H. E. S. S. Collaboration et al. 2015). Considering the presence of multiple bright LMXBs in NGC 4278, we cannot dismiss the possibility that the detection of 1LHAASO J1219+2915 could be attributed to these LMXBs within NGC 4278.
@@ 5.3 NGC 4278: a BL Lac or a LINER
The compact radio morphology and flat radio spectrum of NGC 4278 suggest its similarity to BL Lacs (Giroletti et al. 2005). The SED of NGC 4278 can be represented by a one-zone synchrotron+SSC model, similar to the typical TeV BL Lacs. We compare the derived jet physical parameters of NGC 4278 with that of a TeV-BL Lac sample from Zhang et al. (2012). It seems that NGC 4278 does not fit in with the characteristics of those typical TeV BL Lacs. As shown in Figure 6, both the and values for NGC 4278 are considerably smaller, while is slightly larger compared to typical TeV BL Lacs. In the plane (Figure 7), NGC 4278 well follows the sequence of typical TeV BL Lacs but resides at the low-power end. Nevertheless, when examining the plane, it becomes evident that NGC 4278 stands apart from these TeV BL Lacs due to an unusually smaller ratio of to .
@@
To further investigate this issue, we compile the radio luminosity at 8 GHz () of NGC 4278 and these typical TeV BL Lacs from the Radio Fundamental Catalog (RFC)444https://astrogeo.smce.nasa.gov/rfc/, and plot the -ray luminosity at 1 TeV () against in Figure 8(a). No significant correlation between and for these typical TeV BL Lacs is observed. Moreover, NGC 4278 stands out as it is clearly separated from these typical BL Lacs and positioned at the lower end of the luminosity scale. The luminosity of NGC 4278 is more than two orders of magnitude lower than that of typical TeV BL Lacs in both radio and -ray bands.
@@
As described above, NGC 4278 does exhibit distinctive characteristics that differentiate it from typical TeV BL Lacs, and it is generally classified as a LINER (Heckman 1980) or a ‘bona fide’ CSO (Kiehlmann et al. 2024). In Figure 8(b), we plot the radio luminosity () against the Eddington ratio () for a sample of Seyfert galaxies and LINERs from Sikora et al. (2007), as well as for a sample of BL Lacs from Zhang et al. (2020). Note that the BL Lac sample in Figure 8(b) differs partially from that in Figure 8(a). Not all BL Lacs shown in Figure 8(b) are TeV sources, and not all BL Lacs depicted in Figure 8(a) have available data. However, there is an overlap between the two samples. For Seyfert galaxies and LINERs, represents the luminosity at 5 GHz, while for NGC 4278 and these BL Lacs it corresponds to , where the values are also derived from the observation data obtained in RFC. Both samples (neither considering NGC 4278) show a clear correlation between and . The Pearson correlation analysis yields a coefficient and a chance probability for all Seyfert galaxies and LINERs, while and are obtained for these BL Lacs. In the plane, NGC 4278 is situated within the region occupied by Seyfert galaxies and LINERs, slightly deviating from the low-luminosity extension of the 95% confidence band of the best linear fit for BL Lacs. The Eddington ratio of NGC 4278 is more than one order of magnitude lower than that of BL Lacs, which may be the reason for its low luminosity. Considering all these aspects, NGC 4278 does not resemble a BL Lac; instead, it exhibits more similarities to Seyfert galaxies and LINERs.
@@
NGC 4278 was classified as a LINER-type AGN and no associated -ray source counterpart reported in the 4FGL-DR4 (Abdollahi et al. 2022; Ballet et al. 2023). By reanalyzing all the Fermi-LAT observation data from 2008 Aug 4 to 2024 Apr 24 around the radio position of NGC 4278, no significant -ray emission associated with NGC 4278 is found. Interestingly, NGC 4278 has been detected by LHAASO and is reported to be associated with 1LHAASO J1219+2915 (Cao et al. 2024). To investigate the origin of VHE emission from this source, we examine its X-ray observations obtained by Swift-XRT, Chandra, and XMM-Newton, and find only one valid pointing (on MJD 59546) from Swift-XRT during the LHAASO observations. Surprisingly, the derived X-ray flux from Swift-XRT is more than one order of magnitude higher than the archive X-ray flux obtained by Chandra observation 11.7 years ago, indicating that the X-ray emission of NGC 4278 is currently in a historically high-flux state. Together with the mildly relativistic jets detected in the radio band using the VLBA observations, we propose that the TeV -rays of 1LHAASO J1219+2915 originate from an atypical BL Lac-type AGN in the center of NGC 4278 that has recently become more active; however, other possibilities cannot be completely ruled out yet. The derived parameters for its jet using a one-zone synchrotron+SSC model are also significantly different from that of these typical TeV BL Lacs, with smaller and , but higher energies of electrons. In both radio and TeV bands, NGC 4278 shows luminosity lower by over two orders of magnitude when compared to these typical TeV BL Lacs. In the plane, NGC 4287 slightly deviates from the low-luminosity and low-Eddington-ratio extension of the 95% confidence band representing best linear fit for BL Lacs; instead it exhibits similarities with Seyfert galaxies and LINERs.

### v2

@@ 3 X-ray Emission of the nucleus in NGC 4278
The X-ray light curve of NGC 4278 is compiled in Figure 1, incorporating the archived X-ray data from Younes et al. (2010) and the analysis results from the Chandra and Swift-XRT observations conducted in this study. Notably, NGC 4278 exhibits significant flux variations in the X-ray band. The Swift-XRT observations reveal an exceptionally high-flux state for NGC 4278 in the X-ray band, comparable to the historically highest flux obtained from the XMM-Newton observation on 2004 May 23. The ratio of the flux observed by Swift-XRT in 2021 to that observed by Chandra in 2010 exceeds one order of magnitude. Unfortunately, no additional data between 2010 to 2021 could be acquired to estimate variability timescales accurately. While previous studies have reported a 10% increase in flux over a period of 1 hr during XMM-Newton observations (Younes et al. 2010), due to limited exposure time during the Swift-XRT observations, further investigation into short timescale variability is not feasible in this study.
It should be noted that radio observations at 6 cm from 1972 to 2003 indicate variability in the flux density of NGC 4278, including an observed outburst around 1985, which may be associated with the presence of an active nucleus (Giroletti et al. 2005). The compact and symmetric two-sided radio morphology of NGC 4278 fulfills the criteria for classification as a CSO, and it is even included in a ‘bona fide’ CSO sample (Kiehlmann et al., 2024). CSOs may be transient or episodic sources (O’Dea & Saikia 2021 for a review), with some -ray emitting CSOs reported to display episodic nuclear activity (Lister et al. 2020; Gan et al. 2021, 2022). Recently, the LHAASO collaboration (Cao et al. 2024b) has reported that 1LHAASO J1219+2915 (associated with NGC 4278) exhibits an active period of VHE -ray emission in its 2.5-year light curve, specifically occurring between 2021 Aug 23 (MJD 59449) to 2022 Jan 10 (MJD 59589). It is noteworthy that the data point representing Swift-XRT observation in Figure 1 primarily corresponds to the X-ray flux of NGC 4278 on 2021 Nov 28 (MJD 59546). The observed high-flux state of NGC 4278 by Swift-XRT in 2021 could potentially be attributed to the reactivation of its nucleus, considering the previously low X-ray flux during the Chandra observations in 2010. This may be the reason that NGC 4278 has recently been detected at the VHE -ray band by LHAASO.
The photon spectral index () against the X-ray flux is depicted in Figure 2. Despite the presence of large errors in the data points, the spectral variation can still be observed. However, the X-rays of NGC 4278 do not exhibit the clear behavior of harder when brighter, which is commonly observed among BL Lacs (e.g., Tramacere et al. 2009; Tramacere et al. 2011; Zhang et al. 2013; Di Gesu et al. 2022; Hu et al. 2024).
@@ 4 Constructing and Modeling the SED of Nucleus
The reactivation of the nucleus, as mentioned above, likely results in the detection of VHE -rays and a high-flux state in X-rays from NGC 4278. Consequently, it is plausible that both the emission in X-rays and -rays originate from the same region. In order to further investigate the VHE -ray emission properties of NGC 4278, we collect its multiwavelength observation data from literature and construct a broadband SED. However, only a few non-simultaneous observations from Younes et al. (2010) are available in the radio-optical-UV bands, as shown in Figure 3. The sensitivity curves of LHAASO (Wang et al. 2022) and Fermi-LAT (Funk et al. 2013) are also presented in Figure 3. The hard X-ray spectrum and soft TeV -ray spectrum should correspond to two distinct radiation components, resembling those typical TeV-emitting BL Lacs that have a high synchrotron radiation peak (e.g., Zhang et al. 2012). We employ a simple one-zone leptonic model, widely used in modeling the broadband SEDs of TeV BL Lacs (Tavecchio et al. 2010; Zhang et al. 2012; Yan et al. 2014), to reproduce the SED of NGC 4278, only considering synchrotron and synchrotron-self-Compton (SSC) processes of the relativistic electrons.
@@
The emission region is assumed as a sphere with radius , magnetic field , and a Doppler boosting factor , where the value is fixed at cm, which is a typical value in SED modeling of TeV BL Lacs (Tavecchio et al. 2010; Zhang et al. 2012). The VLBA observations indicate that NGC 4278 has mildly relativistic jets () with a viewing angle () of , which result in a Doppler booting factor of 2.7 (Giroletti et al. 2005). Thus, we first fix to be 2.7. The values of and are constrained by the spectral indices at X-ray and TeV bands. Since cannot be constrained and is fixed as , while is poorly constrained by the last observation point at TeV band. We adjust the parameters , , and to represent the observed SED of NGC 4278. The Klein–Nishina effect and the absorption of extragalactic background light (EBL; Finke et al. 2022) are also taken into account in the SED modeling. The SED fitting result is presented in Figure 3. We obtain mG, , and cm-3 (see also in table 2). The derived value of mG is consistent with the estimated value of mG obtained by assuming that the electrons are cooling through the inverse Compton process on monthly timescale (Cao et al. 2024b).
Considering that the Doppler factor estimated from the radio data may underestimate the one in the -ray emission region of NGC 4278, we also adopt a fixed value of during the SED modeling. This choice is based on previous SED modeling studies for TeV BL Lacs (Tavecchio et al. 2010; Zhang et al. 2012), where almost represents a lower-limit value of these TeV BL Lacs. In this case, a smaller magnetic field strength of mG is obtained. It should be noted that the fitting parameters and are degenerate (Zhang et al. 2012). If a high magnetic field strength is required to represent the broadband SED of NGC 4278, then a smaller . For example, would be necessary with G.
Using the data from the LHAASO first catalog (Cao et al. 2024a), Wang et al. (2024) also considered the one-zone synchrotron+SSC model to explain the X-ray and TeV -ray emission of NGC 4278. They obtained G by assuming and a bulk Lorentz factor of , while for and , they obtained G. It should be noted that their work adopted a significantly smaller value of cm, compared to ours and other BL Lacs in their study. Recently, Dutta & Gupta (2024) utilized the identical dataset and model to fit NGC 4278’s SED as we did, reporting mG with and cm, which aligns consistently with our findings.
As discussed above, the fitting parameters cannot be fully constrained due to the limited observational data of NGC 4278. We only obtain a set of model parameters that can provide an acceptable fit. Based on these SED fitting parameters, we also calculate the jet power () in case of the e± pair jet for NGC 4278 by assuming that the jet consist of electrons (), magnetic fields (), and radiation (). In the case of , the corresponding values are erg s-1, erg s-1, and erg s-1, respectively. On the other hand, when , the corresponding values are erg s-1, erg s-1, and erg s-1. These results are given in Table 2.
@@ 5 Discussion
The angular resolution of LHAASO can achieve around 10 TeV (Ma et al. 2022). As stated in Section 2.1, there are two other GeV–TeV BL Lacs located within a distance of less than from NGC 4278. Moreover, NGC 4278 harbors a substantial number of LMXBs (Brassington et al. 2009; D’Abrusco et al. 2014), and -rays have been detected from several LMXBs in the Galaxy. We are wondering whether the VHE -rays detected by LHAASO from 1LHAASO J1219+2915 could potentially originate from either the two -ray emitting BL Lacs or these LMXBs in NGC 4278.
@@ 5.1 Two TeV BL Lacs Located near NGC 4278
The GeV–TeV emitting BL Lacs 1ES 1218+304 and B2 1215+30 are situated at a distance of 0.929 and 0.964, respectively, from the radio position of NGC 4278. Firstly, we conduct a literature search to identify the highest observed flux at the TeV band for the two BL Lacs. We then extrapolate their intrinsic -ray spectrum up to 100 TeV assuming a power-law spectral form. Subsequently, we calculate the observable spectral shape by taking into account the absorption of EBL and compare it with that of 1LHAASO J1219+2915. Two EBL models are considered in this analysis, corresponding to the strongest absorption (Franceschini & Rodighiero 2017) and weakest absorption (Kneiske & Raue 2010), respectively.
@@
The LHAASO is capable of detecting both 1ES 1218+304 and B2 1215+30 when they are in a high-flux state, as depicted in Figure 4. However, the estimated observable spectra of the two BL Lacs fail to represent the observed highest energy photon of 1LHAASO J1219+2915 even considering the weakest EBL absorption model. In other words, due to their high redshift, it appears impossible for the LHAASO to detect -ray photons with energies exceeding 10 TeV from the two BL Lacs, even when considering the weakest EBL absorption model. Therefore, the VHE -ray emission from 1LHAASO J1219+2915 should not be contributed by the two BL Lacs.
@@ 5.2 LMXBs in NGC 4278
Two identified LMXBs (1SXPS J042749.2-670434 and PSR J1023+0038), along with seven associated LMXBs that are not definitively identified, have been reported in the 4FGL-DR4 (Abdollahi et al. 2022; Ballet et al. 2023). Additionally, several other Galactic LMXBs are also suggested to be the -ray sources (e.g., Xing & Wang 2015; Harvey et al. 2022; Kantzas et al. 2022). In NGC 4278, there exist multiple LMXBs with an X-ray luminosity exceeding erg s-1 (Brassington et al. 2009; Fabbiano et al. 2010), which are likely to be sources of -ray emission as well. Therefore, we estimate the potential -ray flux from these LMXBs in NGC 4278 by assuming that they exhibit a similar X-ray to -ray flux ratio as the Galactic -ray emitting LMXBs, i.e.,
@@
We extrapolate the GeV spectrum of PSR J1023+0038 up to the TeV band by assuming that it follows the same power-law distribution (the black solid line in Figure 5). Multiplying the black solid line by 0.19, we obtain the possible -ray flux (the blue solid line) of the brightest LMXB in NGC 4278. The blue dashed line in Figure 5 presents the result after considering EBL absorption. It is evident that if there is only one -ray emitting LMXB in NGC 4278, its emission would not be detectable by Fermi-LAT but could potentially be detected by LHAASO. NGC 4278 hosts several LMXBs with X-ray luminosity exceeding erg s-1, some of which exhibit significant variability (Fabbiano et al. 2010). Therefore, it is plausible that the observed -rays from 1LHAASO J1219+2915 can be attributed to some LMXBs in NGC 4278. However, no detection has been made for any LMXB at TeV energy band.
In addition, the spectrum at the TeV band of LMXB may not follow the same power-law spectral shape as that at the GeV band; for example, a flatter spectrum is observed in high-mass X-ray binaries (HMXBs) at the TeV band, such as HESS J1018–589 A/1FGL J1018.6–5856 (H. E. S. S. Collaboration et al. 2015). Considering the presence of multiple bright LMXBs in NGC 4278, we cannot dismiss the possibility that the detection of 1LHAASO J1219+2915 could be attributed to these LMXBs within NGC 4278.
@@ 5.3 NGC 4278: a BL Lac or a LINER
The compact radio morphology and flat radio spectrum of NGC 4278 suggest its similarity to BL Lacs (Giroletti et al. 2005). The SED of NGC 4278 can be represented by a one-zone synchrotron+SSC model, similar to the typical TeV BL Lacs. We compare the derived jet physical parameters of NGC 4278 with that of a TeV-BL Lac sample from Tavecchio et al. (2010) and Zhang et al. (2012). The characteristics of NGC 4278 appear to deviate from those typical TeV BL Lacs, as shown in Figure 6. For =10, the value of NGC 4278 is comparable to other BL Lacs, but a smaller value is required. Even when considering a significantly smaller value of =2.7 compared to other BL Lacs, the derived value remains lower than that of most other BL Lacs. In the plane (Figure 7), NGC 4278 follows the sequence of typical TeV BL Lacs but resides at the low-power end. Nevertheless, when examining the plane, it becomes evident that NGC 4278 stands apart from these TeV BL Lacs due to an unusually smaller ratio of to .
@@
The flux densities at 1 TeV of some typical BL Lacs are documented in Zhang et al. (2012). We thus compile the radio luminosity at 8 GHz () of NGC 4278 and these BL Lacs from the Radio Fundamental Catalog (RFC)666https://astrogeo.smce.nasa.gov/rfc/, and plot the -ray luminosity at 1 TeV () against in Figure 8(a). No significant correlation between and for these typical TeV BL Lacs is observed. Moreover, NGC 4278 stands out as it is clearly separated from these typical BL Lacs and positioned at the lower end of the luminosity scale. The luminosity of NGC 4278 is more than two orders of magnitude lower than that of these TeV BL Lacs in both radio and -ray bands.
In Figure 8(b), we plot the radio luminosity () against the Eddington ratio () for a sample of Seyfert galaxies and LINERs from Sikora et al. (2007)777The only exclusion is NGC 1275 due to its classification as a TeV -ray emitting radio galaxy (Aleksić et al. 2012b)., along with a sample of BL Lacs and -ray emitting NLS1s from Zhang et al. (2020) and references therein, as well as two Seyfert galaxies888Only four associated Seyfert galaxies, which have not yet been definitively identified, are presented in the 4FGL-DR4. PKS 0000–160 and NVSS J01531–1844 lack available data and are not included in this study. from the 4FGL-DR4 (Abdollahi et al. 2022; Ballet et al. 2023). The values of for the two Seyfert galaxies (TXS 2116–077 and Circinus galaxy) are taken from Zhang et al. (2020) and Greenhill et al. (2003), respectively. Note that the BL Lac sample in Figure 8(b) differs partially from that in Figure 8(a). Not all BL Lacs shown in Figure 8(b) are TeV sources, and not all BL Lacs depicted in Figure 8(a) have available data. However, there is an overlap between the two samples. For Seyfert galaxies and LINERs from Sikora et al. (2007), represents the luminosity at 5 GHz, while for other sources it corresponds to , where the values are also derived from the observation data in the RFC.
@@
In the plane, NGC 4278 is situated within the region occupied by Seyfert galaxies and LINERs, distinctly separated from these BL Lacs. Circinus galaxy also shares similarities with these Seyfert galaxies and LINERs, while TXS 2116–077 obviously belongs to -ray emitting NLS1s, as suggested in D’Ammando (2019). The BL Lac sample and Seyfert galaxy & LINER sample both (neither including NGC 4278) exhibit a clear correlation between and , whereas this correlation is absent in -ray emitting NLS1s (including TXS 2116–077). The Pearson correlation analysis yields a coefficient of and a chance probability of for the Seyfert galaxy & LINER sample, while obtaining with for these BL Lacs, and deriving with for the -ray emitting NLS1s. The position of NGC 4278 slightly deviates from the low-luminosity extension of the 95% confidence band of the best linear fit for BL Lacs. However, the Eddington ratio of NGC 4278 is more than one order of magnitude lower than that of BL Lacs, which may be the reason for its low luminosity. Considering all these aspects, NGC 4278 does not resemble a BL Lac; instead, it exhibits more similarities to Seyfert galaxies and LINERs.
@@
NGC 4278 was classified as a LINER-type AGN and no associated -ray source counterpart reported in the 4FGL-DR4 (Abdollahi et al. 2022; Ballet et al. 2023). By reanalyzing all the Fermi-LAT observation data from 2008 Aug 4 to 2024 Apr 24 around the radio position of NGC 4278, no significant -ray emission associated with NGC 4278 is found. Interestingly, NGC 4278 has been detected by LHAASO and is reported to be associated with 1LHAASO J1219+2915 at a significance level of 8.8 (Cao et al. 2024b). To investigate the origin of VHE emission from this source, we examine its X-ray observations obtained by Swift-XRT, Chandra, and XMM-Newton, and find only one valid pointing (on 2021 Nov 28, MJD 59546) from Swift-XRT during the LHAASO observations. Moreover, this particular pointing was conducted during an active period (from August 23, 2021 to January 10, 2022) for the source 1LHAASO J1219+2915. Interestingly, the derived X-ray flux from Swift-XRT is more than one order of magnitude higher than the archive X-ray flux obtained by Chandra observation 11.7 years ago, indicating that the X-ray emission of NGC 4278 is in a high-flux state during the active phase of the source 1LHAASO J1219+2915. Together with the mildly relativistic jets detected in the radio band using the VLBA observations, we propose that the TeV -rays of 1LHAASO J1219+2915 originate from an atypical BL Lac-type AGN in the center of NGC 4278 that has recently become more active; however, other possibilities cannot be completely ruled out yet. The derived parameters for its jet using a one-zone synchrotron+SSC model are also significantly different from that of these typical TeV BL Lacs. In both radio and TeV bands, NGC 4278 shows luminosity lower by over two orders of magnitude when compared to these typical TeV BL Lacs. In the plane, NGC 4287 slightly deviates from the low-luminosity and low-Eddington-ratio extension of the 95% confidence band representing best linear fit for BL Lacs; instead it exhibits similarities with Seyfert galaxies and LINERs.